(-14 – 6i) – (28 + 13i) I need to put this in standard form, how do i do it?
Answer by Alan3354(69443)

(-14 – 6i) – (28 + 13i) I need to put this in standard form, how do i do it?
-----------
Add the real parts, and add the imaginary parts.
-14 - 28 = -42
-6i - 13i = -19i
--------
-42 - 19i